Add & Subtract Mixed Number
This a word problem
A women is 64*(3/4)inches tall and her daughter is 61*(11/18)inches tall. How much taller is the women?

Answer by rfer(16322)

64*3/4=48
61*11/18=37.3
48-37.3=10.7in